# NSNQuant: A Double Normalization Approach for Calibration-Free Low-Bit Vector Quantization of KV Cache

### v1

@@ E.4 LM-Eval tasks
For evaluation on GSM8K, HumanEval, CoQA, and MMLU, we use the following tasks provided in lm-evaluation-harness [13]: gsm8k_cot, humaneval, coqa, mmlu_flan_cot_fewshot_humanities, mmlu_flan_cot_fewshot_stem, mmlu_flan_cot_fewshot_social_sciences, mmlu_flan_cot_fewshot_other.
@@ F.1 PPL evaluation in generation setting
Since the perplexity (PPL) evaluation in Table 2 is different from our generation setting, we provide PPL evaluation results in the generation setting. We obtain the output logits by processing tokens one-by-one, just like in a generation scenario. We also adopt residuals to maintain recent tokens in full-precision. The result is presented in Table 15. All methods achieve lower PPLs compared to the results in Table 2 because of the residual, but their relative order has similar trends. Since the PPL evaluation setting in the main table is much more efficient to measure and closer to convention in the previous studies, we use the same setting in the ablation studies.
Table 15: Perplexity evaluated in generation setting with residual, for different models on WikiText-2 and C4.
@@ F.2 Additional results on LongBench
To reduce the impact of noisiness, we evaluate each method in a different perspective: how well each method preserves the original outputs. To acheive this, we measure the ROUGE-L score of each method by comparing their outputs with FP16 outputs. Since Qasper, TREC, and TriviaQA are evaluated using exact matching-based metrics, we exclude them from the task list. The result is shown in Table 17. It clearly shows that NSNQuant is the most effective method which preserves the original output faithfully. On the other hand, CQ is even worse than KIVI + Had, suggesting that calibration-based VQ suffers from significant performance degradation when applied to diverse tasks.
@@
Model Method Bits Qasper QMSum MultiNews TREC TriviaQA SAMSum LCC RepoBench-P Avg.
@@ F.3 Additional results on GSM8K, HumanEval, CoQA, and MMLU
NSNQuant-1bit 1.23 61.64 29.27 60.28 58.73 42.70 59.76 60.94
F.4 Results on AIME-2024
To evaluate the long-context reasoning ability, we evaluate our methods on AIME-2024, with DeepSeek-R1-Distill-Llama-8B. As suggested in DeepSeek [15], we set temperature to 0.6, top-p to 0.95, and max generated tokens to 32768. Since AIME-2024 only contains 30 problems, we run evaluation with 5 different seeds.
The result is shown in Table 19. In 2-bit quantization, NSNQuant and CQ both show little accuracy drop with small differences. On the other hand, under 1-bit quantization, both methods experience significant performance degradation, but NSNQuant-1b achieves over 2$\times$ higher accuracy than CQ-8c10b. While CQ-4b9c is generally strong in math reasoning tasks such as GSM-8K and AIME-2024, CQ-8b10c fails severely in such datasets.
@@ F.4 Results on AIME-2024
F.5 Comparison with QuaRot and SpinQuant
QuaRot [2] and SpinQuant [29] are the recent quantization methods for KV cache which also leverage the Hadamard transform. However, we don’t include them as our baselines because they are not suitable for low-bit quantization. Their evaluation results on GSM8K, HumanEval, CoQA, and MMLU with LLaMA3.1-8B-Instruct are shown in Table 20. We use a clipping ratio of 0.95, following QuaRot, and use a group size of 128 for both keys and values. We only train R2 rotation matrix in SpinQuant, since we don’t apply quantization to weights or activations. We also apply the residual policy of NSNQuant, following the setting in the main experiments. The result shows that QuaRot and SpinQuant underperform in every dataset, and they are worse than our weakest baseline, KIVI-2.
@@ G.1 Inter-channel correlation after the NSN and Hadamard
Figure 10: Layer-wise mean absolute correlation (MAC) between different channels. The results are measured on WikiText2 with LLaMA2-7B.
@@ G.3 Channel-wise mean and standard deviation
Appendix H Latency Measurements
NSNQuant introduces additional computations for normalization, averaging, and shift operations. To assess the overhead, we measure the inference latency under the following setting: a batch size of 32 and an input length of 512. We set the residual size to 64 and generate 64 tokens to ensure that vector quantization is performed exactly once. Latency is measured on a Linux server equipped with a single A100-80GB GPU. Each measurement is repeated 100 times, and the average elapsed time is reported. As shown in Table 21, NSNQuant incurs approximately a 37% increase in inference time during the prefilling stage. In contrast, decoding is significantly faster, as NSNQuant substantially reduces the overhead caused by intensive memory access.
Table 21: Inference latency comparison between the baseline and NSNQuant. We report the average elapsed time over 100 runs on a single A100-80GB GPU. While NSNQuant introduces moderate overhead during the prefilling stage due to additional computations, it achieves faster decoding by alleviating memory access bottlenecks.
Method Prefill Decode
FP16 1.229s 0.0507s
NSNQuant-2b 1.675s 0.0324s
NSNQuant-1b 1.698s 0.0324s
Appendix I Comparison with CQ
@@
Although CQ and NSNQuant both quantize multiple channels jointly, their motivations behind adopting vector quantization (VQ) differ. CQ employs VQ to capture correlations across channels, whereas NSNQuant is inspired by the observation from QuIP# [37] that VQ performs particularly well when quantizing ball-shaped vectors. Since the channel distribution is aligned with the standard normal distribution regardless of model and data, we can build an optimized codebook for compressing the normal distribution data in advance.
@@
Appendix J Related Work
Recent efforts to reduce the memory and latency bottlenecks of LLMs have largely focused on weight quantization and KV cache compression. Weight quantization methods such as GPTQ [12] and AWQ [27] significantly reduce the memory footprint of model weights while preserving accuracy. In particular, vector quantization (VQ) methods like QuIP#[37], AQLM[11], and VPTQ [28] enable extremely low-bit quantization of model weights. However, these approaches do not mitigate the growing memory consumption of the KV cache, which scales linearly with sequence length and model depth during long-context inference.
To address this, recent studies have explored compressing the KV cache. KVQuant [19] and KIVI [30] observe the presence of outlier patterns in KV representations and propose to apply channel-wise quantization for key and token-wise quantization for value. CQ [43] extends non-uniform quantization (nuq) to multiple channels, achieving effective 1-bit KV cache compression. Orthogonal approaches, such as H2O [44] and SnapKV [25], reduce memory through cache eviction of uninformative tokens. Hybrid strategies like ZipCache [17] and MiKV [41] selectively maintain high-precision cache for important tokens while applying low-bit quantization elsewhere. Our method uses the same precision for all tokens, but future work could incorporate token importance as in hybrid approaches.
Appendix K Limitation and broader impacts
K.1 Limitation

### v2

@@ E.4 LM-Eval tasks
For evaluation on GSM8K, HumanEval, CoQA, and MMLU, we use the following tasks provided in lm-evaluation-harness [eval-harness]: gsm8k_cot, humaneval, coqa, mmlu_flan_cot_fewshot_humanities, mmlu_flan_cot_fewshot_stem, mmlu_flan_cot_fewshot_social_sciences, mmlu_flan_cot_fewshot_other.
@@ F.1 PPL evaluation in generation setting
Since the perplexity (PPL) evaluation in Table 2 is different from our generation setting, we provide PPL evaluation results in the generation setting. We obtain the output logits by processing tokens one-by-one, just like in a generation scenario. We also adopt residuals to maintain recent tokens in full-precision. The result is presented in Table 17. All methods achieve lower PPLs compared to the results in Table 2 because of the residual, but their relative order exhibits similar trends. Since the PPL evaluation setting in the main table is much more efficient to measure and closer to convention in the previous studies, we use the same setting in the ablation studies.
Table 17: Perplexity measured in the generation setting with residual. We evaluate LLaMA2-7B and LLaMA3.1-8B on WikiText-2 and C4.
@@ F.2 Additional results on LongBench
To reduce the impact of noisiness, we evaluate each method from a different perspective: how well each method preserves the original outputs. To achieve this, we measure the ROUGE-L score of each method by comparing their outputs with FP16 outputs. Since Qasper, TREC, and TriviaQA are evaluated using exact matching-based metrics, we exclude them from the task list. The result is shown in Table 19. It clearly shows that NSNQuant is the most effective method which preserves the original output faithfully. On the other hand, CQ is even worse than KIVI + Had, suggesting that calibration-based VQ suffers from significant performance degradation when applied to diverse tasks.
@@
Model Method Bits
Qasper
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RepoBench-P
Avg.
@@ F.3 Additional results on GSM8K, HumanEval, CoQA, and MMLU
NSNQuant-1b 1.23 61.64 29.27 60.28 58.73 42.70 59.76 60.94
F.4 Additional latency breakdown results
Table 21 presents additional latency breakdown results under different configurations, measured with LLaMA2-7B. We fix the number of generated tokens to 64 and vary the batch size and input prompt length to observe performance trends. The results exhibit a similar pattern to Table 5, where NSNQuant incurs additional overhead during the prefill stage but demonstrates lower latency during the decode stage.
Table 21: Latency (ms) breakdown of NSNQuant-2b compared with the FP16 baseline. BS, L, HT and VQ refer to batch size, input prompt length, Hadamard transform and vector quantization, respectively. The number of generated tokens is fixed to 64.
BS L Stage Method Total NSN HT VQ
4 4096 Prefill FP16 1294.84
NSNQuant-2b 1761.45 104.13 14.69 324.16
Decode
(per step)
FP16 49.34
NSNQuant-2b 33.17 0.29 1.70 0.14
32 512 Prefill FP16 1225.90
NSNQuant-2b 1707.22 155.22 15.04 326.85
Decode
(per step)
FP16 50.80
NSNQuant-2b 36.49 0.34 1.88 0.66
128 128 Prefill FP16 1223.20
NSNQuant-2b 1698.50 105.76 15.02 329.62
Decode
(per step)
FP16 63.45
NSNQuant-2b 45.25 0.83 1.87 2.43
F.5 Results on AIME-2024
To evaluate the long-context reasoning ability, we evaluate our methods on AIME-2024, with DeepSeek-R1-Distill-Llama-8B. As suggested in DeepSeek [guo2025deepseek], we set temperature to 0.6, top-p to 0.95, and maximum generated tokens to 32768. Since AIME-2024 only contains 30 problems, we run the evaluation with 5 different seeds.
[…]
@@ F.5 Results on AIME-2024
F.6 Comparison with QuaRot and SpinQuant
QuaRot [ashkboos2025quarot] and SpinQuant [liu2024spinquant] are the recent quantization methods for KV cache which also leverage the Hadamard transform. However, we don’t include them as our baselines because they are not suitable for low-bit quantization. Their evaluation results on GSM8K, HumanEval, CoQA, and MMLU with LLaMA3.1-8B-Instruct are shown in Table 23. We use a clipping ratio of 0.95, following QuaRot, and use a group size of 128 for both keys and values. We only train R2 rotation matrix in SpinQuant, since we don’t apply quantization to weights or activations. We also apply the residual policy of NSNQuant, following the setting in the main experiments. The result shows that QuaRot and SpinQuant underperform in every dataset, and they are worse than our weakest baseline, KIVI-2.
@@ G.1 Inter-channel correlation after the NSN and Hadamard
Figure 10: Layer-wise mean absolute correlation (MAC) between different channels. The results are measured on WikiText-2 with LLaMA2-7B.
@@ G.3 Channel-wise mean and standard deviation
Appendix H Comparison with CQ
@@
Although CQ and NSNQuant both quantize multiple channels jointly, their motivations behind adopting vector quantization (VQ) differ. CQ employs VQ to capture correlations across channels, whereas NSNQuant is inspired by the observation from QuIP# [pmlr-v235-tseng24a] that VQ performs particularly well when quantizing ball-shaped vectors. Since the channel distribution is aligned with the standard normal distribution regardless of model and data, we can build an optimized codebook for compressing the normal distribution data in advance.
@@
Appendix I Limitation and broader impacts
I.1 Limitation